Question 11 of 14 Map pling r the following reacti
Answer #1

E cell = 1.92 - 2 * 0.575 = 0.77

we know ecell = e(not) cell - o.o591/n log k

e cell =0 , due to equlibrium

k = 0.777 * 2 /o.o591 = 10^2.62